Circular No. 8928

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V2468 CYGNI = NOVA CYGNI 2008 N. N. Samus, Institute of Astronomy, Russian Academy of Sciences,

informs us that this nova (cf. IAUC 8927) has been assigned the variablestar designation V2468 Cyg.

M. M. Beaky, Truman State University, reports that a CCD spectrum (range 450–850 nm, resolution 0.4 nm/pixel) of V2468 Cyg, acquired with the 0.36-m telescope at the Truman Observatory on Mar. 11.46 UT, shows Hα (FWHM = 790 km/s) and Hβ lines to be prominent in emission, and numerous Fe II emission lines are also visible. Thus, the nova belongs to the ‘Fe II’ class of Williams (1992, A.J. 104, 725).

CCD photometry of V2468 Cyg by K. Hornoch, Ondrejov, Czech Republic, with a 0.65-m reflector: Mar. 10.175 UT, V = 8.27 ± 0.01, B–V = +1.10 ± 0.03, V–I = +1.52 ± 0.02 (comparison star HIP 98323). Visual magnitude estimates, in part from E. O. Waagen, AAVSO: Mar. 9.417, 7.7 (J. E. Bortle, Stormville, NY, U.S.A.); 10.187, 8.3 (Hornoch); 12.362, 8.1 (M. Bradbury, Greenwood, IN, U.S.A.); 13.452, 9.2 (L. Shotter, Uniontown, PA, U.S.A.); 15.020, 9.0 (B. Granslo, Fjellhamar, Norway); 17.443, 8.6 (Shotter); 19.056, 9.6 (Granslo).

(3749) BALAM F. Marchis, Carl Sagan Center (CSC), SETI Institute; J. Pollock, Ap-

palachian State University (ASU); P. Pravec, Ondřejov Observatory; M. Baek, CSC; J. Greene and L. Hutton, ASU; P. Descamps, Institut de Mécanique Céleste et de Calcul des Ephémérides; and D. E. Reichart, K. M. Ivarsen, J. A. Crain, M. C. Nysewander, A. P. Lacluyze, J. B. Haislip, and J. S. Harvey, University of North Carolina at Chapel Hill, report that photometric observations obtained during 2007 July 15 to Oct. 19 reveal that the central object of (3749) is actually a binary system with a mutual orbital period of 33.38 ± 0.02 hr. The primary shows a rotation period of 2.80483 ± 0.00002 hr and a lightcurve amplitude of 0.13 mag. Mutual eclipse/occultation events that were 0.15 and 0.43-mag deep on July 18 and Sept. 8, respectively, indicate a size ratio of the two inner components of 0.4 ± 0.1. The system’s parameters need to be refined with further observations. A distant satellite of (3749) was discovered in 2002 (see IAUC 7827) at a distance of 0′′.47 using adaptive optics (AO) imaging capabilities. A separation between the two inner components showing mutual events is estimated to be roughly 20 km, corresponding to an angular separation of about 0′′.02, not yet attainable with current AO systems.

2008 March 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green